Give exact and approximate solutions to 3 decimal places.
(x - 4)^2 = 81
The exact solutions are: __________
The approximate solutions are: ____________

Answer by rapaljer(4671)

(x - 4)^2 = 81
Square root each side:

x-4= +9
x=13

x-4=-9
x-4+4=-9+4
x=-5

R^2